### v1

@@ IV-A Functional Form for the Terrain Model
{dmath}
f(x^j,y^j) = ∑_n=1^N a_ncos(ω_1,n x^j +ω_2,n y^j ) + b_nsin(ω_3,n x^j +ω_4,n y^j ),
@@ IV-B Differentiable Wheel-Terrain Interaction
$g_{i}(\mathbf{x}_{k},\mathbf{u}_{k})=0,$ (25)
@@
$\mathbf{u}_{k}^{*}(\mathbf{x}_{k})=\min_{\mathbf{u}_{k}}\left\lVert g_{i}(\mathbf{x}_{k},\mathbf{u}_{k})\right\rVert_{2}^{2}.$ (26)
@@
Consider the NLS problem (26). We can define $\mathbf{H}=\mathrm{D}_{\mathbf{u}_{k}\mathbf{u}_{k}}^{2}\mathbf{g}(\mathbf{x}_{k},\mathbf{u}_{k}(\mathbf{x}_{k}))\in\mathbb{R}^{m\times m}$, where $\mathbf{g(.)}$ is obtained by stacking $g_{i}(.)$ and $\mathbf{B}=\mathrm{D}_{\mathbf{x_{k}u_{k}}}^{2}\mathbf{g}(\mathbf{x}_{k},\mathbf{u}_{k}(\mathbf{x}_{k}))\in\mathbb{R}^{m\times n}$. Then, the Jacobian of the optimal pose is
@@ IV-C Bi-Level Optimization Based Trajectory Planning
$\mathbf{u}_{k}^{*}(\mathbf{x}_{k})=\min_{\mathbf{u}_{k}}\left\lVert g_{i}(\mathbf{x}_{k},\mathbf{u}_{k})\right\rVert_{2}^{2}.$ (28f)
@@ V-B Validation using the Gazebo simulator
In this section, we analyze the accuracy of our pose predictor by comparing the ground truth data obtained from Gazebo with the values obtained by solving the NLS problem (26) with the help of (17). To this end, we drive the Husky robot in Gazebo on synthetic terrain models and collect the odometry data, as shown in Fig. 4(a). We then compare the ground truth values $z_{gt}$, $\beta_{gt}$, and $\gamma_{gt}$ from Gazebo with the predicted $z_{pred}$, $\beta_{pred}$, and $\gamma_{pred}$ from our pose optimizer. Several runs of the robot on different terrains were performed, and some qualitative results are shown in Figures 4(b) and 4(c). We can see that the predicted pose values closely match the ground truth values. The statistical results presented in Figures 4(d) and 4(e) show the minimum, maximum, and median error obtained across several runs. It can be seen that our NLS pose optimizer is highly accurate in predicting the pose of the robot on different terrains, and the error values are minimal.
@@ V-C Validation of the Safe Planning on Uneven Terrain
Next, we present a statistical result in Fig. 7(a). Here, we also highlight the significance of the weighting term $w_{\theta}$ in the stability cost 16. Several trajectories were simulated on different terrains, and the average worst-case tip-over angles were recorded. It can be seen that for the trajectories without stability cost, the value is lower. In the case of tip-over angle, higher positive values are preferable, as shown in (13). The values with stability cost are far higher than those without stability cost, which confirms that the trajectories taken by the vehicle with stability cost have a higher safety margin. It can also be seen that as the weight $w_{\theta}$ increased from 0.05 to 0.2, the values of the angles increased, which in turn increased the tip-over stability margin.

### v2

@@ IV-A Functional Form for the Terrain Model
$f(x^{j},y^{j})=\sum_{n=1}^{N}a_{n}\cos{(\omega_{1,n}x^{j}+\omega_{2,n}y^{j})}+b_{n}\sin{(\omega_{3,n}x^{j}+\omega_{4,n}y^{j})},$ (20)
@@ IV-B Differentiable Wheel-Terrain Interaction
$g_{j}(\mathbf{x}_{k},\mathbf{u}_{k})=0,\quad j=1\dots 16,$ (31)
@@
$\mathbf{u}_{k}^{*}(\mathbf{x}_{k})=\operatorname*{arg\,min}_{\mathbf{u}_{k}}\sum_{j}g_{j}(\mathbf{x}_{k},\mathbf{u}_{k})^{2}.$ (32)
@@
Consider the NLS problem (32). We can define $\mathbf{H}=\mathrm{D}_{\mathbf{u}_{k}\mathbf{u}_{k}}^{2}\mathbf{g}(\mathbf{x}_{k},\mathbf{u}_{k}(\mathbf{x}_{k}))\in\mathbb{R}^{m\times m}$, where $\mathbf{g(.)}$ is obtained by stacking $g_{j}(.)$ and $\mathbf{B}=\mathrm{D}_{\mathbf{x_{k}u_{k}}}^{2}\mathbf{g}(\mathbf{x}_{k},\mathbf{u}_{k}(\mathbf{x}_{k}))\in\mathbb{R}^{m\times n}$. Then, the Jacobian of the optimal pose is
@@ IV-C Bi-Level Optimization Based Trajectory Planning
$\mathbf{u}_{k}^{*}(\mathbf{x}_{k})=\operatorname*{arg\,min}_{\mathbf{u}_{k}}\sum_{j}g_{j}(\mathbf{x}_{k},\mathbf{u}_{k})^{2}.$ (34f)
@@ V-B Validation using the Gazebo simulator
In this section, we analyze the accuracy of our pose predictor by comparing the ground truth data obtained from Gazebo with the values obtained by solving the NLS problem (32) with the help of (19). To this end, we drive the Husky robot in Gazebo on synthetic terrain models and collect the odometry data, as shown in Fig. 3(a). We then compare the ground truth values $z_{gt}$, $\beta_{gt}$, and $\gamma_{gt}$ from Gazebo with the predicted $z_{pred}$, $\beta_{pred}$, and $\gamma_{pred}$ from our pose optimizer. Several runs of the robot on different terrains were performed, and some qualitative results are shown in Figures 3(b) and 3(c). We can see that the predicted pose values closely match the ground truth values. The statistical results presented in Figures 3(d) and 3(e) show the minimum, maximum, and median error obtained across several runs. It can be seen that our NLS pose optimizer is highly accurate in predicting the pose of the robot on different terrains, and the error values are minimal.
@@ V-C Validation of the Safe Planning on Uneven Terrain
Next, we present a statistical result in Fig. 6(a). Here, we also highlight the significance of the weighting term $w_{\theta}$ in the stability cost 18. Several trajectories were simulated on different terrains, and the average worst-case tip-over angles were recorded. It can be seen that for the trajectories without stability cost, the value is lower. In the case of tip-over angle, higher positive values are preferable, as shown in (15). The values with stability cost are far higher than those without stability cost, which confirms that the trajectories taken by the vehicle with stability cost have a higher safety margin. It can also be seen that as the weight $w_{\theta}$ increased from 0.05 to 0.2, the values of the angles increased, which in turn increased the tip-over stability margin.